### v4

# MVTamperBench: Evaluating Robustness of Vision-Language Models

Abstract
Multimodal Large Language Models (MLLMs), also known as Large Multi-modal Models (LMMs), are recent advancement of Vision-Language Models (VLMs), that have driven major advances in video understanding, yet their vulnerability to adversarial tampering and manipulations remains underexplored. To address this gap, we introduce MVTamperBench, a benchmark that systematically evaluates MLLM robustness against five prevalent tampering techniques: rotation, masking, substitution, repetition, and dropping. Built from 3.4K original videos—expanded to over 17K tampered clips spanning 19 video tasks.
MVTamperBench challenges models to detect manipulations in spatial and temporal coherence. We evaluate 45 recent MLLMs from 15+ model families, revealing substantial variability in resilience across tampering types and showing that larger parameter counts do not necessarily guarantee robustness. MVTamperBench sets a new benchmark for developing tamper-resilient MLLM in safety-critical applications, including detecting clickbait, preventing harmful content distribution, and enforcing policies on media platforms. We release all code [^1] https://amitbcp.github.io/MVTamperBench/ and data [^2] https://huggingface.co/datasets/Srikant86/MVTamperBench to foster open research in trustworthy video understanding.
MVTamperBench: Evaluating Robustness of Vision-Language Models
Amit Agarwal1*, Srikant Panda2*, Angeline Charles3*, Bhargava Kumar4, Hitesh Patel5, Priyaranjan Pattnayak6, Taki Hasan Rafi7, Tejaswini Kumar 4, Dong-Kyu Chae7+ 1Liverpool John Moores University 2Birla Institute of Technology 3Christ University 4Columbia University 5New York University 6University of Washington 7Hanyang University *Equal Contribution +Leadership Correspondence: amit.pinaki@gmail.com, dongkyu@hanyang.ac.kr
1 Introduction
Multimodal Large Language Models (MLLMs) have driven significant advancements in video understanding, enabling applications across domains such as surveillance, healthcare, and autonomous systems. However, their increasing adoption on high-usage platforms (e.g., Instagram, Facebook) has exposed critical vulnerabilities. Tampered videos are frequently weaponized to evade platform policies, spread harmful content, and propagate clickbait, posing significant challenges for content moderation and policy enforcement Kingra et al. (2023); Times of India (2024). These vulnerabilities underscore the urgent need to enhance the robustness of MLLMs against real-world manipulations.
While adversarial robustness has been extensively studied in static image contexts, video tampering presents unique challenges. Techniques such as frame dropping, masking, repetition, substitution, and rotation disrupt both spatial and temporal coherence, often causing severe model failures. Existing adversarial approaches, including black-box Jiang et al. (2019) and cross-modal Wei et al. (2021) attacks, primarily address isolated scenarios rather than systematically evaluating diverse tampering types.
Existing multimodal benchmarks Pattnayak et al. (2024); AGARWAL (2021)—such as MMBench-Video Fang et al. (2024), BLINK Fu et al. (2025), and Video-MME Fu et al. (2024)—focus on multimodal comprehension and temporal reasoning but overlook adversarial robustness. For example, MMBench-Video evaluates cross-modal alignment without adversarial testing; BLINK targets long-form temporal reasoning without addressing tampering impacts; and Video-MME, while effective for vision-language alignment, omits tampering-specific tasks. This leaves a critical gap in systematically assessing how MLLMs withstand real-world manipulations.
To bridge this gap, we introduce MVTamperBench, a benchmark specifically designed to evaluate MLLM robustness against five prevalent tampering techniques. Built from 3.4K original videos—expanded into over 17K tampered clips spanning 19 video tasks—MVTamperBench challenges models to detect manipulations by distrupting temporal and spatial coherence in a video.
Our main contributions are as follows:
  • We introduce MVTamperBench, a benchmark that systematically evaluates MLLMs on five major video manipulations, focusing on spatial and temporal coherence to stimulate real-world scenarios.
  • We propose a unified evaluation methodology that frames tampering detection as a multiple-choice task, enabling straightforward, interpretable and consistent performance comparisons.
  • Through experiments on 45 MLLMs across 15+ families, we identify critical vulnerabilities across MLLM families, without any correlation between model size and performance.
By focusing on tampering resilience, MVTamperBench provides a critical tool for improving MLLM robustness. It enables the development of tamper-resilient models applicable to real-world challenges such as clickbait detection, content moderation, and policy enforcement, advancing adversarial robustness in video understanding.
2 Related Work
The development of benchmarks for MLLMs has significantly advanced the evaluation of image and video understanding tasks, encompassing spatial reasoning, temporal comprehension, object detection, and common sense inference. However, the robustness of MLLMs to adversarial manipulations, particularly video tampering, remains underexplored. This section reviews existing benchmarks, categorizing them into image-based and video-based evaluations, and critically analyzes their contributions and shared limitations.
2.1 Image-Based Understanding
Benchmarks like BLINK Fu et al. (2025) and MuirBench Wang et al. (2024a) focus on evaluating static visual reasoning. BLINK tests foundational tasks such as depth estimation, forensic detection, and visual correspondence, providing a diverse set of challenges for spatial reasoning. Similarly, MuirBench extends this evaluation to multi-image tasks, including action recognition and geographic reasoning, by synthesizing information from diverse sources. While these benchmarks have advanced static image understanding, their reliance on single or multiple still images excludes temporal dynamics and limits their applicability to scenarios involving sequential manipulations, such as those found in video content.
2.2 Video-Based Understanding
Video-based benchmarks have expanded the scope of evaluation by incorporating temporal and multimodal reasoning. MVBench Li et al. (2024e), MMBench-Video Fang et al. (2024), and Video-MME Fu et al. (2024) focus on tasks such as event detection, episodic reasoning, and contextual understanding. These benchmarks challenge models with diverse tasks spanning object interactions, long-duration video analysis, and domain-specific reasoning. Similarly, LongVU Shen et al. (2024) introduces spatio-temporal compression techniques to enhance efficiency, while MotionEpic Fei et al. (2024a) integrates Spatial-Temporal Scene Graphs (STSG) for fine-grained cognitive tasks. However, while these benchmarks assess temporal coherence and reasoning, they do not systematically address adversarial robustness, particularly in tampering scenarios.
Other benchmarks, such as Wolf Li et al. (2024a) and Sharingan Chen et al. (2024a), target specialized video understanding tasks. Wolf focuses on captioning using a mixture-of-experts strategy, while Sharingan extracts action sequences from desktop recordings using frame-differential approaches. Although these benchmarks achieve high accuracy in their domains, they are limited to specific tasks and lack general mechanisms to evaluate the resilience to tampering.
2.3 Identified Gaps and Motivation
Despite their contributions to advancing MLLM evaluation, existing benchmarks largely focus on performance under ideal conditions, neglecting robustness to tampering or adversarial effects. Techniques such as frame substitution, masking, repetition, dropping, and rotation disrupt temporal coherence and pose unique challenges for multimodal models. The absence of systematic evaluations of these tampering techniques highlights a critical gap in ensuring model reliability for real-world applications like forensic analysis, media verification, and misinformation detection.
Figure 1: Illustration of Video Frame Tampering Techniques. Each row shows a specific tamper type applied to a video snippet, highlighting the respective impact on temporal/spatial coherence.
To address these challenges, we introduce MVTamperBench, a comprehensive benchmark designed to systematically evaluate MLLMs robustness against video tampering effects. By introducing diverse manipulations, MVTamperBench broadens the evaluation landscape, enabling a deeper understanding of model strengths and vulnerabilities under adversarial scenarios.
3 MVTamperBench
We introduce MVTamperBench, our benchmark specifically designed to evaluate the robustness of MLLMs against diverse video tampering techniques. In the following subsections, we detail its construction, design choices, and key features.
3.1 Benchmark Construction
To evaluate MLLM robustness under adversarial conditions, we apply distinct tampering methods—Dropping, Masking, Substitution, Repetition, and Rotation—to the 3,487 original MVBench Li et al. (2024e) videos (excluding NTU dataset due to licensing), resulting in a total of 17,435 tampered clips. These manipulations target both spatial and temporal coherence, simulating common real-world tampering scenarios such as deliberate frame editing or slicing from unrelated content.
3.2 Tampering Techniques
Dropping: Removes a 1-second segment, creating temporal discontinuity.
Masking: Overlays a black rectangle on a 1-second segment, simulating visual data loss.
Rotation: Rotates a 1-second segment by 180 degrees, introducing spatial distortion.
Substitution: Replaces a 1-second segment with a pre-selected clip from another video, disrupting temporal and contextual flow.
Repetition: Repeats a 1-second segment, introducing temporal redundancy.
By applying these effects uniformly across all videos, MVTamperBench ensures consistent and comparable evaluation.
3.3 Design & Implementation
Tampering Duration (1 Second). We fix tampering to a 1-second segment to align with reports of minimal yet impactful real-world tampering, e.g., short edits in social media or subtle modifications in surveillance feeds Alicia (2024). Preliminary experiments revealed that using less than 1 second could be overlooked by certain model sampling mechanisms, whereas longer tampering (>3s) often resembled normal scene transitions, reducing adversarial impact.
Tampering Location (Middle). All manipulations occur at the video’s midpoint to disrupt central content. Pilot tests showed that tampering near the start or end risked mimicking scene cuts or information loss, making detection less indicative of genuine adversarial robustness.
Substitution Source. For Substitution, the 1-second clip is randomly chosen from a consistent pool of different videos within MVBench. This ensures uniform difficulty across all samples, preventing confounds from domain shifts or overly simplistic substitutes.
Modular & Scalable Framework. Each technique is encapsulated in a reusable class, facilitating custom parameterization (e.g., tampering duration, location, intensity). Our open-source code integrates seamlessly with VLMEvalKit Duan et al. (2024), promoting reproducible experiments and easy extension to other datasets or tasks.
3.4 Dataset Scope and Statistics
All 3,487 MVBench videos undergo each of the five tampering types, producing 17,435 tampered clips. Figure 2 illustrates the distribution of their durations, capturing diverse scenarios from short (3-5s) to extended (>20s) sequences. By uniformly applying manipulations, MVTamperBench ensures comparability across different MLLM architectures and training regimes.
Figure 2: The dataset spans a broad range of durations, reflecting varied real-world conditions.
3.5 Summary
By enforcing consistent parameters (1s duration, midpoint placement) and systematically applying five tampering methods, MVTamperBench offers a controlled yet flexible platform for evaluating tampering resilience. Our design can be easily extended with additional manipulations (e.g., noise injection, partial masking) or integrated into domain-specific contexts like surveillance feeds analysis or clickbait detection. We further compare MVTamperBench with other benchmarks, details on source of videos and tasks in Appendix 7.1
Figure 3: F1 scores across models and tampering types. High-performing models are robust across all types.
4 Experimentation and Results
This section outlines the experimental setup, results, and analysis of 45 models evaluated on the benchmark, highlighting their strengths, weaknesses, and actionable insights. We provide further overview of the evaluated MLLMs in Appendix 7.3.
Evaluation Protocol Each model is tasked with identifying whether a video has been tampered with or not, simulating real-world scenarios where detecting manipulations is critical. For every video, the model is presented with the following prompt:
Does this video exhibit any signs of tampering, such as corruption, blackouts, rotated frames, repeated frames, or swapped frames?
Options: A. Yes     B. No
The dataset comprises both tampered and untampered videos. For each tampered video, the corresponding untampered video is included, ensuring a balanced distribution of tampered and untampered samples. Models must select one of the two options (Yes or No), and their predictions are compared against ground truth labels to determine correctness. This task is repeated for all tampering types, enabling a comprehensive assessment of model robustness across diverse manipulations.
Metrics The primary evaluation metric is the F1 Score, chosen for its ability to balance precision and recall, particularly in scenarios where misclassifications (false positives and false negatives) can significantly impact robustness evaluation. This is particularly important in our setup, where models must not only detect manipulations but also avoid false positives on non-tampered videos. To capture performance across all tampering types, we compute individual F1 scores for each tampering type. The F1 (overall) score is then calculated as the macro average of these individual F1 scores, ensuring all tampering types contribute equally to the final evaluation. This approach effectively highlights per-tamper-type strengths and weaknesses while providing a holistic measure of overall model performance.
Figure 4: Distribution of F1 (overall) scores across models.
Figure 5: Correlation between model size and F1 (overall), with a Pearson value of 0.05.
4.1 Results and Analysis
We evaluate the performance of 45 MLLMs using F1 (overall) scores and individual F1 scores across five tampering types. The distribution of F1 (overall) scores (Figure 4) reveals significant variability in model robustness, with several models struggling to detect tampering (F1 < 0.2), while a few high-performing models achieve F1 > 0.8.
We observe no correlation between model size and tampering detection performance (Pearson correlation = 0.05, Figure 5). This highlights that architectural differences and training techniques, rather than parameter count, contribute more significantly to robustness against tampering.
Adaption to Tampering Techniques
Figure 3 showcases model-specific adaptability across different tampering types. Models with F1 (overall) > 0.8 are consistently robust across all tampering types, while those with F1 < 0.2 perform slightly better on Masking, which relies more on spatial reasoning. Models in between generally struggle with temporal disruptions like Dropping and Substitution, reflecting challenges in temporal coherence.
4.1.1 Models Analysis Based on Performance
Figure 6 categorizes MLLMs into low-, moderate-, and high -performing groups based on their F1-score distribution. We define the boundaries using the 0.25 quantile (F1 = 0.071) and 0.75 quantile (F1 = 0.846). To round-off the quantiles thresholds to the nearest integer, making the low-performing models (F1 (overall) < 0.01) and high-performing models (F1 (overall) > 0.8).
Figure 7 illustrates the average F1 scores for each category, revealing stark differences between groups. High-performing models maintain consistent performance across all tampering types, whereas low-performing models show significant weaknesses, particularly for tampering effects that disrupt temporal coherence (Dropping, Substitution). Moderate-performing models excel slightly in Masking (Figure 8).
Figure 6: Distribution of models into low, moderate, and high-performing categories based on F1 (overall) scores across tampering types.
Figure 7: Average F1 (overall) scores across low, moderate, and high-performing model categories.
Figure 8: F1 (overall) scores for tampering types across model categories. Low & Moderate-performing models perform slightly better on Masking, while high-performing models show consistent robustness.
Figure 9: Variance in F1 (overall) scores across tampering types for each model category.
We analyze variance in model performances across tampering types in Figure 9. High-performing models exhibit negligible variance, indicating consistent robustness across all tampering effects. This stability suggests that these models are well-equipped to handle both spatial & temporal disruptions introduced by tampering.
Low-performing models, display minimal variance across most tampering types, with the exception of Masking, which shows a slightly higher variance. This highlights a specific weakness in processing visual obfuscations, likely due to their reliance on static features rather than robust temporal reasoning. The minimal variance across other tampering types suggests that these models fail uniformly, regardless of the type of manipulation.
Moderate-performing models demonstrate the highest variance overall, particularly for tampering types such as Dropping, Repetition, and Substitution. This behavior indicates an inconsistency in their ability to adapt to tampering effects. While these models achieve a balanced performance across easier tampering types, temporal disruptions like Dropping, Repetition, and Substitution pose significant challenges, leading to occasional spikes in variance. This suggests that moderate-performing models, while more capable than low-performing counterparts, still struggle to maintain robustness across a diverse set of tampering scenarios. We further analyze individual model performances in Appendix 7.3.1
4.1.2 Models Analysis Based on Model Size
We categorize models based on their parameter sizes into small (<7B), medium (7B–26B), and large (>26B) groups (Figure 11). While Figure 12 shows that larger models generally achieve higher F1 (overall) scores, Figure 5 confirms no significant correlation between model size and tampering detection performance (Pearson correlation = 0.05). A closer examination of individual model trends across size categories reveals several noteworthy patterns discussed in Appendix 7.3.2
Figure 10: F1 (overall) scores for task categories in MVTamperBench. Tasks like Episodic Reasoning achieve higher scores, while Counterfactual Inference is more challenging.
Trends Across Families.
Across size categories, we observe distinct trends in model performance. The VILA model family consistently improves with size, with the exception of VILA1.5-8B, highlighting the scalability of its architecture for tampering detection. Similarly, Qwen2-VL demonstrates significant gains with increased parameters, though it trails behind other families in absolute performance.
The Llama3.2-Vision family, despite its scaling efforts, reveals the diminishing returns of increasing model size without architectural or training advancements. Meanwhile, Molmo models illustrate the importance of efficient design, as Molmo-1B outperforms its larger variants like Molmo-72B. Finally, the dominance of InternVL-2.5 across all sizes highlights the benefits of balanced architecture & task-specific training strategies.
4.1.3 Analysis Across Video Task Types
MVTamperBench comprises 19 video task categories, each evaluated across five tampering techniques. Figure 10 highlights the F1 (overall) scores for each task, averaged across all tampering types and models. While certain tasks exhibit high detection F1 score, others remain significantly more challenging. Appendix 7.3.3 provides additional insights into the performance trends of each tampering type across different tasks.
Figure 11: Distribution of Models by Size; Categories: Small (<7B), Medium (7B–26B), and Large (>26B)
Figure 12: F1 (overall) scores for models by size category. Larger models generally tend to perform better.
Easier Tasks.
Tasks such as Episodic Reasoning, Scene Transition, Ego-centric Navigation, and State Change consistently achieve higher F1 scores, as shown in Figure 10. These tasks often involve shorter temporal dependencies and simpler spatial reasoning, allowing models to rely on pre-trained vision-language features rather than advanced temporal reasoning.
For example, in Episodic Reasoning, tampering effects like Dropping and Repetition minimally affect task performance, as models primarily focus on memorizing what it has seen. Similarly, Scene Transitions tasks depend more on detecting static or localized changes, making them less susceptible to temporal disruptions such as Dropping or Repetition. The results indicate that models trained on extensive image-based datasets excel in tasks with lower temporal complexity.
Challenging Tasks.
Tasks such as Action Prediction, Counterfactual Inference, and Fine-Grained Action pose significant challenges for tampering detection. These tasks inherently require complex temporal reasoning and context preservation, making them highly sensitive to disruptions introduced by tampering techniques.
In Counterfactual Inference, tampering effects like Substitution & Dropping cause significant confusion, as they disrupt the continuity of the video narrative required for hypothetical reasoning. Similarly, Fine-Grained Action detection is heavily impacted by Rotation, distorting spatial relationships crucial for identifying subtle movements.
Interestingly, Action Prediction tasks also highlight the limitations of current MLLMs in understanding and reasoning about temporal progression. Models often fail to account for disruptions in video sequences, leading to degraded performance across tampering types.
We further discuss the Key Findings & Future Directions in Appendix 7.5 & 7.6 respectively.
5 Conclusion
We introduced MVTamperBench, a comprehensive benchmark for evaluating the robustness of Multimodal Large Language Models (MLLMs) against five key video tampering techniques—Dropping, Masking, Repetition, Rotation, and Substitution. Through systematic experiments on 19 video tasks involving 45 models across 15+ families, we observed pronounced variability in resilience. Notably, even LMMs exceeding 70B parameters suffer severe performance drops, whereas select small models (<7absent7<7< 7B) demonstrate unexpectedly strong tampering detection, illustrating that size alone does not ensure robustness.
Beyond revealing these vulnerabilities, MVTamperBench offers actionable insights for refining model architectures and training pipelines, highlighting gaps that must be addressed before these systems can be reliably deployed in safety-critical settings. Our open-source framework supports reproducible evaluations and community-driven extensions, enabling researchers to integrate additional datasets or adapt tampering methods for domains like clickbait detection, content moderation, and surveillance feeds analysis.
Looking ahead, we plan to expand MVTamperBench with new tampering types (e.g., noise injection, frame shuffling) and domain-specific scenarios (e.g., healthcare, surveillance). By illuminating the nuanced impacts of video manipulations and guiding innovation in robust LMM architectures, MVTamperBench lays a strong foundation for next-generation multimodal models capable of withstanding adversarial manipulations.
6 Limitation
MVTamperBench provides a robust framework for evaluating LMM resilience against video tampering, but there are limitations that present opportunities for future work.
First, while the benchmark evaluates five tampering techniques, expanding to additional manipulations to better capture subtle, and emerging techniques. Second,the dataset currently relies on limited datasets. Expanding the benchmark to incorporate videos from diverse sources, such as user-generated content, surveillance footage, or policy enforcement, would broaden its applicability and relevance to task-specific and domain-specific challenges. Third, while the current evaluation focuses on binary detection, future benchmarks could assess a model’s ability to classify the specific tampering type, providing deeper insights into its robustness. Finally, scalability to closed-source and extremely large-scale models (>100absent100>100> 100B parameters) remains a challenge due to computational and cost constraints.
Addressing these limitations will enable MVTamperBench to further advance tampering detection and resilience across a broader range of applications.
Acknowledgement
This work was partly supported by (1) the National Research Foundation of Korea(NRF) grant funded by the Korea government(MSIT)(RS-2024-00345398) and (2) the Institute of Information & communications Technology Planning & Evaluation(IITP) grant funded by the Korea government(MSIT)(RS-2020-II201373, Artificial Intelligence Graduate School Program (Hanyang University)).