### v5

# MVTamperBench: Evaluating Robustness of Vision-Language Models

Abstract
Multimodal Large Language Models (MLLMs), are recent advancement of Vision-Language Models (VLMs) that have driven major advances in video understanding. However, their vulnerability to adversarial tampering and manipulations remains underexplored. To address this gap, we introduce MVTamperBench, a benchmark that systematically evaluates MLLM robustness against five prevalent tampering techniques: rotation, masking, substitution, repetition, and dropping; based on real-world visual tampering scenarios such as surveillance interference, social media content edits, and misinformation injection. MVTamperBench comprises  3.4K original videos, expanded into over  17K tampered clips covering 19 distinct video manipulation tasks. This benchmark challenges models to detect manipulations in spatial and temporal coherence. We evaluate 45 recent MLLMs from 15+ model families. We reveal substantial variability in resilience across tampering types and show that larger parameter counts do not necessarily guarantee robustness. MVTamperBench sets a new benchmark for developing tamper-resilient MLLM in safety-critical applications, including detecting clickbait, preventing harmful content distribution, and enforcing policies on media platforms. We release all code, data, and benchmark to foster open research in trustworthy video understanding.
MVTamperBench: Evaluating Robustness of Vision-Language Models
Amit Agarwal1*+, Srikant Panda2*, Angeline Charles3*, Hitesh Laxmichand Patel5, Bhargava Kumar4, Priyaranjan Pattnayak6, Taki Hasan Rafi7, Tejaswini Kumar4, Hansa Meghwani1, Karan Gupta5, Dong-Kyu Chae7+ 1Liverpool John Moores University 2Birla Institute of Technology 3Christ University 4Columbia University 5New York University 6University of Washington 7Hanyang University *Equal Contribution +Corresponding Authors Correspondence: amit.pinaki@gmail.com, dongkyu@hanyang.ac.kr
1 Introduction
Multimodal Large Language Models (MLLMs) have catalyzed significant progress in video understanding, enabling a wide array of applications across domains such as surveillance, healthcare, and autonomous systems. However, their growing integration into high-traffic platforms (e.g., Instagram, Facebook, TikTok) has exposed critical vulnerabilities. Specifically, tampered videos are increasingly exploited to bypass platform policies, disseminate harmful content, and promote clickbait, posing serious challenges for content moderation and policy enforcement Kingra et al. (2023); Times of India (2024).
These threats underscore the urgent need to improve the robustness of MLLMs against real-world manipulations. Although video tampering can involve audio modifications, synthetic speech overlays, or deepfake generation, our benchmark focuses exclusively on visual-only manipulations. This choice is driven by the current limitations of existing models, most of which currently lack comprehensive audiovisual processing capabilities.
Unlike adversarial robustness in static images, an area that has been extensively studied, video tampering introduces unique challenges that arise from the interplay of spatial and temporal dynamics. Common manipulation techniques such as frame dropping, masking, repetition, substitution, and rotation disrupt this coherence, frequently resulting in catastrophic model failures. These methods mirror real-world adversarial tactics: substitution injects objectionable material (e.g., nudity, violence) to circumvent detection; dropping eliminates key surveillance evidence; repetition loops footage to obscure illicit activity; masking occludes critical regions; and rotation induces spatial distortions commonly seen in edited or re-uploaded content.
Existing adversarial and traditional approaches, including black-box Jiang et al. (2019) and cross-modal Wei et al. (2021) attacks, primarily address isolated scenarios rather than systematically evaluating diverse tampering types. Furthermore, existing multimodal benchmarks Pattnayak et al. (2024); Agarwal (2021)—such as MMBench-Video Fang et al. (2024), BLINK Fu et al. (2025), and Video-MME Fu et al. (2024)—focus on multimodal comprehension and temporal reasoning but overlook adversarial robustness. For example, MMBench-Video evaluates cross-modal alignment without adversarial testing; BLINK targets long-form temporal reasoning without addressing tampering impacts; and Video-MME, while effective for vision-language alignment, omits tampering-specific tasks. This leaves a critical gap in systematically assessing how MLLMs withstand real-world manipulations.
To bridge this gap, we introduce MVTamperBench, a benchmark specifically designed to evaluate MLLM robustness against five prevalent tampering techniques. Built from 3.4K original videos—expanded into over 17K tampered clips spanning 19 video tasks—MVTamperBench challenges models to detect manipulations by disrupting temporal and spatial coherence in a video. By focusing on tampering resilience, we believe that MVTamperBench provides a critical tool for improving MLLM robustness, and eventually, it enables the development of tamper-resilient models applicable to real-world challenges such as clickbait detection, content moderation, and policy enforcement, advancing adversarial robustness in video understanding for high-stakes domains.
Our main contributions are as follows:
  • We introduce MVTamperBench, a benchmark that systematically evaluates MLLMs on five major video manipulations, focusing on spatial and temporal coherence to stimulate real-world scenarios.
  • We propose a unified evaluation methodology that frames tampering detection as a multiple-choice task, enabling straightforward, interpretable and consistent performance comparisons.
  • Through experiments on 45 MLLMs across 15+ families, we identify critical vulnerabilities across MLLM families. across MLLM families, without any correlation between model size and performance.
  • Our released code enables researchers to integrate additional datasets and adapt or add new tampering, facilitating domain-specific extensions, and supports reproducibility.
2 Related Work
The development of benchmarks for MLLMs has significantly advanced the evaluation of image and video understanding tasks. They have covered spatial reasoning, temporal comprehension, object detection, common sense inference, and so on. However, the robustness of MLLMs to adversarial manipulations, particularly video tampering, remains underexplored. This section reviews existing benchmarks, which can be categorized into image-based and video-based evaluations, and analyzes their contributions and shared limitations.
2.1 Image-based Understanding
Benchmarks like BLINK Fu et al. (2025) and MuirBench Wang et al. (2024a) focus on evaluating static visual reasoning. BLINK tests foundational tasks such as depth estimation, forensic detection, and visual correspondence, which cover a diverse set of challenges for spatial reasoning. Similarly, MuirBench extends this evaluation to multi-image tasks, including action recognition and geographic reasoning, by synthesizing information from diverse sources. While these benchmarks have advanced static image understanding, their reliance on single or multiple still images excludes temporal dynamics and limits their applicability to scenarios involving sequential manipulations, such as those found in video content.
2.2 Video-based Understanding
Video-based benchmarks have expanded the scope of evaluation by incorporating temporal and multimodal reasoning. MVBench Li et al. (2024e), MMBench-Video Fang et al. (2024), and Video-MME Fu et al. (2024) focus on tasks such as event detection, episodic reasoning, and contextual understanding. These benchmarks challenge models with diverse tasks spanning object interactions, long-duration video analysis, and domain-specific reasoning. Similarly, LongVU Shen et al. (2024) introduces spatio-temporal compression techniques to enhance efficiency, while MotionEpic Fei et al. (2024a) integrates Spatial-Temporal Scene Graphs (STSG) for fine-grained cognitive tasks. However, while these benchmarks assess temporal coherence and reasoning, they do not systematically address adversarial robustness, particularly in tampering scenarios.
Figure 1: Illustration of the five video frame tampering techniques. Each row shows a specific tamper type applied to a video snippet, describing the respective impact on temporal/spatial coherence.
Other benchmarks, such as Wolf Li et al. (2024a) and Sharingan Chen et al. (2024a), target specialized video understanding tasks. Wolf focuses on captioning using a mixture-of-experts strategy, while Sharingan extracts action sequences from desktop recordings using frame-differential approaches. Although these benchmarks achieve high accuracy in their domains, they are limited to specific tasks and lack general mechanisms to evaluate the resilience to tampering.
In summary, despite their contributions to advancing MLLM evaluation, existing benchmarks largely focus on performance under ideal conditions and neglect robustness to tampering or adversarial effects. Techniques such as frame substitution, masking, repetition, dropping, and rotation disrupt temporal coherence and pose unique challenges for multimodal models. The absence of systematic evaluations of these tampering techniques highlights a critical gap in ensuring model reliability for real-world applications like forensic analysis, media verification, and misinformation detection.
3 MVTamperBench
We introduce MVTamperBench, a comprehensive benchmark designed to systematically evaluate the robustness of MLLMs against video tampering techniques. Through our MVTamperBench which introduces diverse manipulations, we aim to broaden the evaluation landscape, enabling a deeper understanding of model strengths and vulnerabilities under adversarial scenarios. Table 3 compares MVTamperBench with existing benchmarks, highlighting its focus areas, strengths, and unique contributions.
In the following subsections, we detail its construction, design choices, and key features. More details can be found in our code [^1] https://amitbcp.github.io/MVTamperBench/, data [^2] https://hf.co/datasets/Srikant86/MVTamperBench and benchmark  [^3] https://github.com/open-compass/VLMEvalKit repositories.
3.1 Benchmark Construction
To evaluate MLLM robustness under adversarial conditions, we apply distinct tampering methods as shown in Figure 1Dropping, Masking, Substitution, Repetition, and Rotation—to the 3,487 original MVBench Li et al. (2024e) videos (excluding NTU dataset due to licensing), resulting in a total of 17,435 tampered clips. These manipulations target both spatial and temporal coherence, thereby simulating common real-world tampering scenarios such as deliberate frame editing or slicing from unrelated content.
3.2 Tampering Techniques
Dropping: Removes a 1-second segment for creating temporal discontinuity.
Masking: Overlays a black rectangle on a 1-second segment. It aims to simulate visual data loss.
Rotation: Rotates a 1-second segment by 180 degrees for introducing spatial distortion.
Substitution: Replaces a 1-second segment with a pre-selected clip from another video, in order to disrupt temporal and contextual flow.
Repetition: Repeats a 1-second segment, introducing temporal redundancy.
The aforementioned effects are applied uniformly across all videos to ensure consistent and comparable evaluation.
3.3 Design & Implementation
Tampering Duration (1 Second).
We fix tampering to a 1-second segment to align with reports of minimal but impactful real-world tampering, e.g., short edits on social networks or subtle modifications on surveillance feeds Alicia (2024). Our preliminary experiments revealed that using less than 1 second could be overlooked by certain model sampling mechanisms, whereas longer tampering (>3s) often resembled normal scene transitions, reducing adversarial impact.
Tampering Location (Middle).
All manipulations occur at the video’s midpoint to disrupt central content. Our pilot tests showed that tampering near the start or end risked mimicking scene cuts or information loss, which makes detection less indicative of genuine adversarial robustness.
Substitution Source.
For Substitution, the 1-second clip is randomly chosen from a consistent pool of different videos within MVBench. This ensures uniform difficulty across all samples, thus preventing confounds from domain shifts or overly simplistic substitutes.
Modular & Scalable Framework.
Each technique is encapsulated in a reusable class for facilitating custom parameterization (e.g., tampering duration, location, intensity). Our open-source code will integrate seamlessly with VLMEvalKit Duan et al. (2024), thereby promoting reproducible experiments and easy extension to other datasets or tasks.
To inform our final design decisions, we conducted a series of exploratory experiments examining the effects of tampering position and duration. Results from these alternative design configurations are detailed in Appendix A.2.1.
3.4 Dataset Scope and Statistics
All 3,487 MVBench videos undergo each of the five tampering types, producing 17,435 tampered clips. The five manipulations are uniformly applied to ensure comparability across different MLLM architectures and training regimes. Figure 2 illustrates the distribution of their durations. We can observed that diverse scenarios from short (3-5s) to extended (>20s) sequences are included.
Figure 2: Distribution of video durations. Our dataset spans a broad range of durations, which can reflect varied real-world conditions.
3.5 Summary
By enforcing consistent parameters (1s duration, midpoint placement) and systematically applying five tampering methods, MVTamperBench offers a controlled yet flexible platform for evaluating tampering resilience. Our design can be easily extended with additional manipulations (e.g., noise injection, partial masking, positions), deepfakes or integrated into domain-specific contexts like surveillance feeds analysis or clickbait detection. We provide additional details on MVTamperBench, including video sources and associated tasks, in Appendix A.1.
Figure 3: F1 scores across models and tampering types. High-performing models are robust across all types.
4 Experiments
This section outlines the experimental setup, results, and analysis of 45 models evaluated on our proposed benchmark. Our results highlight their strengths, weaknesses, and actionable insights. We provide further overview of the evaluated MLLMs in Appendix A.3.
4.1 Experimental Setups
Evaluation Protocol
Each model is tasked with identifying whether a video has been tampered with or not. For every video, the model is presented with the following structured prompt:
Does this video exhibit any signs of tampering, such as corruption, blackouts, rotated frames, repeated frames, or swapped frames?
Options: A. Yes     B. No
The dataset comprises both tampered and non-tampered videos. For each tampered video, the corresponding non-tampered video is included to ensure a balanced distribution. Models must select one of the two options (Yes or No), and their predictions are compared against ground truth labels to determine correctness. This task is repeated for all tampering types.
We also compared structured, general, and chain-of-thought prompts (Appendix A.2.2), finding that general and CoT variants led to higher false positives due to limited temporal reasoning in current MLLMs.
Metrics
The primary evaluation metric is the F1 Score, chosen for its ability to balance precision and recall, particularly in scenarios where misclassifications (false positives and false negatives) can significantly impact robustness evaluation. We believe that this metric is the most suitable for our setup, where models must not only detect manipulations but also avoid false positives on non-tampered videos. To capture performance across all tampering types, we compute individual F1 scores for each tampering type. In addition, we also measure F1 (overall) score as the macro average of these individual F1 scores. We thus highlight per-tamper-type strengths and weaknesses as well as overall model performance.
Figure 4: Distribution of F1 (overall) scores across models.
Figure 5: Scatter plot showing no correlation between model size and overall F1 (Pearson r=0.05).
4.2 Results and Analysis
We evaluate the performance of 45 MLLMs using F1 (overall) scores and individual F1 scores across five tampering types. The distribution of F1 (overall) scores (Figure 4) reveals significant variability in the robustness of the model, with several models struggling to detect tampering (F1 < 0.2), while a few high-performing models achieve F1 > 0.8.
We did not observe a correlation between model size and tampering detection performance (Pearson correlation = 0.05, Figure 5). This highlights that architectural differences and training techniques, rather than parameter count, contribute more significantly to robustness against tampering.
In addition, Figure 3 showcases model-specific adaptability across different tampering types. Models with F1 (overall) > 0.8 are consistently robust across all tampering types, while those with F1 < 0.2 perform slightly better on Masking, which relies more on spatial reasoning. Models in between generally struggle with temporal disruptions like Dropping and Substitution, reflecting challenges in temporal coherence.
4.2.1 Analysis based on Performance
Figure 6 categorizes MLLMs into low-, moderate-, and high- performing groups based on their F1-score distribution. We define the boundaries using the 0.25 quantile (F1 = 0.071) and 0.75 quantile (F1 = 0.846). We round-off the quantile thresholds to the nearest integers to identify low-performing models (F1 (overall) < 0.01) and high-performing models (F1 (overall) > 0.8).
Figure 7 illustrates the average F1 scores for each category. We observe stark differences between groups: High-performing models maintain consistent performance across all tampering types, whereas low-performing models show significant weaknesses, particularly for tampering effects that disrupt temporal coherence (Dropping, Substitution). Moderate-performing models excel slightly in Masking (Figure 8).
Figure 6: Distribution of Number of Models in low, moderate, and high-performing categories based on F1 (overall).
Figure 7: Average F1 (overall) scores across low, moderate, and high-performing model categories.
Figure 8: F1 scores for tampering types across model categories. Low & Moderate-performing models perform slightly better on Masking, while high-performing models show consistent robustness.
Figure 9: Variance in F1 scores across tampering types for each model category.
We analyze the variance in model performance between tampering types in Figure 9. High-performing models exhibit negligible variance, indicating consistent robustness across all tampering effects. This stability suggests that these models are well-equipped to handle both spatial & temporal disruptions introduced by tampering.
Low-performing models display minimal variance across most tampering types, with the exception of Masking, which shows a slightly higher variance. This highlights a specific weakness in processing visual obfuscations, likely due to their reliance on static features rather than robust temporal reasoning. The minimal variance across other tampering types suggests that these models fail uniformly, regardless of the type of manipulation.
Moderate-performing models demonstrate the highest variance overall, particularly for tampering types such as Dropping, Repetition, and Substitution. This behavior indicates an inconsistency in their ability to adapt to tampering effects. While these models achieve a balanced performance across easier tampering types, temporal disruptions like Dropping, Repetition, and Substitution pose significant challenges, leading to occasional spikes in variance. This suggests that moderate-performing models, while more capable than low-performing counterparts, still struggle to maintain robustness across a diverse set of tampering scenarios. We provide a detailed analysis of individual model performance in Appendix A.4.1, and present qualitative trends based on model architecture and training paradigms across all models studied in Appendix A.4.2, offering insights into the factors influencing performance.
4.2.2 Analysis based on Model Size
We categorize models based on their parameter sizes into small (<7B), medium (7B–26B), and large (>26B) groups (Figure 11). While Figure 12 shows that larger models generally achieve higher F1 (overall) scores, Figure 5 confirms no significant correlation between model size and tampering detection performance (Pearson correlation = 0.05). A closer examination of individual model trends across size categories reveals several noteworthy patterns discussed in Appendix A.4.3.
Figure 10: F1 (overall) scores for task categories in MVTamperBench. Tasks like Episodic Reasoning achieve higher scores, while Counterfactual Inference is more challenging.
Trends Across Families.
Across size categories, we observe distinct trends in model performance. The VILA model family consistently improves with size, with the exception of VILA1.5-8B, highlighting the scalability of its architecture for tampering detection. Similarly, Qwen2-VL demonstrates significant gains with increased parameters, though it trails behind other families in absolute performance.
The Llama3.2-Vision family, despite its scaling efforts, reveals the diminishing returns of increasing model size without architectural or training advancements. Meanwhile, Molmo models illustrate the importance of efficient design, as Molmo-1B outperforms its larger variants like Molmo-72B. Finally, the dominance of InternVL-2.5 across all sizes highlights the benefits of balanced architecture & task-specific training strategies.
4.2.3 Analysis across Video Task Types
MVTamperBench comprises 19 video task categories, each evaluated across five tampering techniques. Figure 10 highlights the F1 (overall) scores for each task, averaged across all tampering types and models. While certain tasks exhibit high detection F1 score, others remain significantly more challenging. Appendix A.4.4 provides additional insights into the performance trends of each tampering type across different tasks.
Figure 11: Distribution of models grouped by size; Categories: Small (<7B), Medium (7B–26B), and Large (>26B).
Figure 12: F1 (overall) scores for models by size category. Larger models generally tend to perform better.
Easier Tasks.
Tasks such as Episodic Reasoning, Scene Transition, Ego-centric Navigation, and State Change consistently achieve higher F1 scores, as shown in Figure 10. These tasks often involve shorter temporal dependencies and simpler spatial reasoning, allowing models to rely on pre-trained vision-language features rather than advanced temporal reasoning.
For example, in Episodic Reasoning, tampering effects like Dropping and Repetition minimally affect task performance, as models primarily focus on memorizing what it has seen. Similarly, Scene Transitions tasks depend more on detecting static or localized changes, making them less susceptible to temporal disruptions such as Dropping or Repetition. The results indicate that models trained on extensive image-based datasets excel in tasks with lower temporal complexity.
Challenging Tasks.
Tasks such as Action Prediction, Counterfactual Inference, and Fine-Grained Action pose significant challenges for tampering detection. These tasks inherently require complex temporal reasoning and context preservation, making them highly sensitive to disruptions introduced by tampering techniques.
In Counterfactual Inference, tampering effects like Substitution & Dropping cause significant confusion, as they disrupt the continuity of the video narrative required for hypothetical reasoning. Similarly, Fine-Grained Action detection is heavily impacted by Rotation, distorting spatial relationships crucial for identifying subtle movements.
Interestingly, Action Prediction tasks also highlight the limitations of current MLLMs in understanding and reasoning about temporal progression. Models often fail to account for disruptions in video sequences, leading to degraded performance across tampering types.
We further discuss our Key Findings & Future Directions in Appendix A.6 & A.7 respectively.
5 Concluding Remarks
Novel Benchmark.
We introduced MVTamperBench, a comprehensive benchmark for evaluating the robustness of Multimodal Large Language Models (MLLMs) against five key video tampering techniques—Dropping, Masking, Repetition, Rotation, and Substitution. Through systematic experiments on 19 video tasks involving 45 models across 15+ families, we observed pronounced variability in resilience. Notably, even MLLMs exceeding 70B parameters suffer severe performance drops, whereas select small models (<7absent7<7< 7B) demonstrate unexpectedly strong tampering detection, illustrating that size alone does not ensure robustness.
Impact on Research Community.
Beyond revealing these vulnerabilities, we believe that MVTamperBench offers actionable insights for refining model architectures and training pipelines, highlighting gaps that must be addressed before these systems can be reliably deployed in safety-critical settings. Our open-source framework will support reproducible evaluations and community-driven extensions, enabling researchers to integrate additional datasets or adapt tampering methods for domains like clickbait detection, content moderation, and surveillance feeds analysis.
Future Work.
Looking ahead, we plan to expand MVTamperBench with new tampering types (e.g., noise injection, frame shuffling) and domain-specific scenarios (e.g., healthcare, surveillance). By illuminating the nuanced impacts of video manipulations and guiding innovation in robust MLLM architectures, MVTamperBench lays a strong foundation for next-generation multimodal models capable of withstanding adversarial manipulations.
6 Limitation
MVTamperBench provides a robust framework for evaluating MLLM resilience against video tampering, but there are limitations that present opportunities for future work.
First, while the benchmark evaluates five tampering techniques, expanding to additional manipulations/modalities to better capture subtle, and emerging techniques. Second,the dataset currently relies on limited datasets. Expanding the benchmark to incorporate videos from diverse sources, such as user-generated content, surveillance footage, or policy enforcement, would broaden its applicability and relevance to task-specific and domain-specific challenges. Third, while the current evaluation focuses on binary detection, future benchmarks could assess a model’s ability to classify the specific tampering type, providing deeper insights into its robustness. Finally, scalability to closed-source and extremely large-scale models (>100absent100>100> 100B parameters) remains a challenge due to computational and cost constraints.
Addressing these limitations will enable MVTamperBench to further advance tampering detection and resilience in a broader range of applications.
Acknowledgement
This work was partly supported by (1) the National Research Foundation of Korea(NRF) grant funded by the Korea government(MSIT)(RS-2024-00345398) and (2) the Institute of Information & communications Technology Planning & Evaluation(IITP) grant funded by the Korea government(MSIT)(RS-2020-II201373, Artificial Intelligence Graduate School Program (Hanyang University)).
Appendix A Appendix
A.1 MVTamperBench Details
Dataset Name Primary Scene Type and Unique Characteristics
STARWu et al. (2024a) Indoor actions and object interactions
PAXIONWang et al. (2023) Real-world scenes with nuanced actions
Moments in Time (MiT) V1Monfort et al. (2019) Indoor/outdoor scenes across varied contexts
FunQAXie et al. (2025) Humor-focused, creative, real-world events
CLEVRERMao et al. (2022) Simulated scenes for object movement and reasoning
Perception TestPatraucean et al. (2024) First/third-person views for object tracking
Charades-STAfor AI (2024) Indoor human actions and interactions
MoVQAZhang et al. (2023) Diverse scenes for scene transition comprehension
VLN-CEKrantz (2024) Indoor navigation from agent perspective
TVQALei et al. (2018) TV show scenes for episodic reasoning
Table 1: Summary of Datasets in MVTamperBench
Table 1 summarizes the datasets included in MVTamperBench, each contributing unique characteristics for robust tampering detection evaluation. We eliminate videos from the NTU dataset in MVTamperBench to avoid conflicting terms in the license. The dataset is then systematically expanded through the application of tampering effects to generate a comprehensive benchmark.
Table 3 summarizes MVTamperBench’s unique contributions compared to existing benchmarks Hu et al. (2024); Song et al. (2024a); Li et al. (2024e); Fu et al. (2024); Xia et al. (2024); Fei et al. (2024b) and use-cases which are spread across videos,images and documents Agarwal et al. (2024a, b, c, d, b, 2025a); Patel et al. (2024); Agarwal and Pachauri (2023). It also compares MVTamperBench with existing benchmarks, highlighting its focus areas, strengths, and unique contributions. We the growing interest area in MLLMS, new benchmarks that require special attention include Svbench Yang et al. (2025), Video-MMLU Song et al. (2025), SEA-VL Cahyawijaya et al. (2025), MotionBench Hong et al. (2025) and synthetic data generation techinques Dua et al. (2024, 2025); Pabolu et al. (2024a, b).
A.2 Ablation Studies
We conduct ablation studies on key design decisions and prompt-engineering strategies to identify optimal configurations for constructing our benchmark. These experiments evaluate the effects of tampering duration, tampering position, and prompt formulation on model performance.
A.2.1 Ablation: Tampering Position and Duration
To understand the influence of tampering characteristics on model performance, we conducted two ablation studies: (1) varying the duration of tampered segments, and (2) altering their position within the video timeline.
Tampering Position Ablation.
We varied the tampering position to occur after 25%, 50%, or 75% of the video timeline (Table 2). Detection performance remained largely stable across these conditions, indicating that most models are relatively insensitive to the specific location of tampering within the video. This suggests a consistent ability to maintain temporal coherence understanding regardless of when the manipulation occurs. We exclude tampering at the very beginning or end of videos, as these segments often coincide with natural scene transitions, delayed starts, or abrupt endings, which could introduce confounding noise into the benchmark.
Category Model 25% 50% 75%
Low Qwen2VL-7B 0.009 0.009 0.009
LLaVaVideo-7B 0.006 0.006 0.006
LLaVaOV-72B 0.0441 0.044 0.044
Moderate Aria 0.7213 0.721 0.721
Qwen2VL-72B 0.352 0.352 0.352
Phi3.5Vision 0.707 0.707 0.707
High VILA1.5-40B 0.879 0.879 0.879
InternVL2.5-8B 0.721 0.7210 0.721
Table 2: Impact of tampering position on F1 scores across models with varying robustness levels. Results are reported for tampering introduced at 25%, 50%, and 75% positions within the video timeline.
Benchmark Scope (Image/Video) Focus Strengths Unique Contributions BLINK Image Visual reasoning Tests spatial relationships in images. Introduces a framework for spatial reasoning with auxiliary sketches. MUIRBENCH Image Multi-image understanding Evaluates complex real-world scenarios. Includes multi-image relations like narrative and complementary. MVBench Video Temporal reasoning, event recognition Assesses video understanding over time. Focuses on video dynamics with temporal task coverage. MMBench-Video Video Long-form video understanding Handles multi-step event recognition in long videos. Evaluates LVLMs on free-form QA using temporal reasoning. Video-MME Video Multi-modal video understanding Evaluates tasks like action recognition and captioning. Combines multiple modalities for enhanced contextual understanding. LongVU Video Spatiotemporal compression Efficiently processes long videos with adaptive compression. Novel spatiotemporal compression mechanism using cross-modal query. MotionEpic Video Object tracking Tracks object interactions across video frames. Implements fine-grained spatial-temporal scene graph reasoning. Wolf Video Video captioning Improves video captioning with expert strategies. Introduces CapScore for LLM-based caption evaluation. Sharingan Video Action sequence extraction Focuses on action recognition in desktop recordings. Proposes differential and direct frame-based methods for user action extraction. Video-of-Thought Video Step-by-step video reasoning Excels in human-like video reasoning with chain-of-thought processes. Integrates spatial-temporal scene graphs (STSG) for fine-grained reasoning. CARES Video Scene comprehension and emotional analysis Analyzes multi-modal emotional and contextual nuances. Integrates context-aware emotional reasoning for enhanced video understanding. Visual-Sketchpad Image Visual interaction and sketching tasks Supports creative and interactive visual reasoning. Bridges sketch-based reasoning with image analysis for enhanced user interaction. MovieChat Video Conversational video understanding Enhances video understanding with conversational context. Introduces dialogue-based comprehension for temporal and narrative reasoning. MVTamperBench (Proposed) Video Tampered video detection Robustly identifies tampered regions in video datasets. Unique focus on domain-specific tampering scenarios with real-world applicability.
Table 3: Enhanced comparison between Video and Image Analysis Benchmarks, with unique contributions highlighted.
Category Model 1s 2s 3s
Low Qwen2-VL-7B 0.009 0.201 0.257
LLaVa-Video-7B 0.006 0.158 0.216
LLaVa-
Onevision-72B 0.044 0.257 0.249
Moderate Aria 0.721 0.779 0.782
Qwen2-VL-72B 0.352 0.485 0.499
Phi3.5-Vision 0.707 0.763 0.769
High VILA1.5-40B 0.879 0.898 0.901
InternVL2.5-8B 0.721 0.801 0.805
Table 4: Impact of tampering duration on F1 scores across models of varying robustness. Results are reported for tampering durations of 1s, 2s, and 3s.
Tampering Duration Ablation.
We evaluated three tampering durations: 1s, 2s, and 3s. Results in Table 4 indicate a consistent improvement in detection performance with increased tampering duration across all models. Lower-performing models (e.g., Qwen2-VL-7B, LLaVa-Video-7B) benefited more significantly, suggesting a dependency on longer anomalous intervals for effective detection. In contrast, top-performing models (e.g., VILA-1.5-40B, InternVL-2.5-8B) exhibited performance saturation, implying diminishing returns beyond a certain duration threshold.
A.2.2 Ablation: Prompt Engineering
To determine the most effective prompt formulation for evaluating MLLM robustness in video tampering detection, we conducted a series of prompt-engineering experiments. Each variant prompts the model to assess whether the video has been tampered with, as aligned with the objective of our benchmark.
  • Structured Prompt (used in benchmark): Explicitly lists tampering types.
    Prompt: Does this video exhibit any signs of tampering, such as corruption, blackouts, rotated frames, repeated frames, or swapped frames?
  • Generic Prompt: Uses general phrasing that mimics natural user queries.
    Prompt: Does this video exhibit any signs of tampering, manipulation, or inconsistency?
  • Chain-of-Thought (CoT) Prompt: Instructs the model to reason step by step before deciding.
    Prompt:
    Scan the video step by step:
    1. 1.
      Check each segment for visual glitches, repeated or missing content, or rotations.
    2. 2.
      Check if any frame seems inconsistent with the rest of the video.
    3. 3.
      Decide if any part looks manipulated or tampered.
    Does the video show any signs of tampering, manipulation, or inconsistency?
Category Model Prompt Type F1 Score
Low Qwen2VL-7B Structured 0.009
Generic 0.001
CoT 0.001
LLaVaVideo-7B Structured 0.006
Generic 0.002
CoT 0.001
LLaVaOV-72B Structured 0.044
Generic 0.001
CoT 0.001
Moderate Aria Structured 0.721
Generic 0.151
CoT 0.148
Qwen2VL-72B Structured 0.352
Generic 0.009
CoT 0.010
Phi3.5Vision Structured 0.707
Generic 0.201
CoT 0.208
High VILA1.5-40B Structured 0.879
Generic 0.458
CoT 0.466
InternVL2.5-8B Structured 0.721
Generic 0.386
CoT 0.389
Table 5: F1 scores of models evaluated with three prompt types: Structured, Generic, and Chain-of-Thought (CoT). Structured prompts consistently yield the highest performance across all robustness categories.
Findings: Structured prompts consistently outperformed both generic and CoT prompts across all model tiers (Table 5). Low- and moderate-performing models exhibited significant performance drops with more open-ended prompts, likely due to limited temporal reasoning capabilities and lack of explicit training. Generic prompts particularly led to frequent false positives, as models misinterpreted benign variability as tampering. These findings support the use of structured prompts to ensure consistent and interpretable evaluation.
Figure 13: Taxonomy of Multimodal Large Language Models (MLLMs), organized by family, version, and first-generation releases.
A.3 Overview of Multimodal Large Language Models (MLLMs)
Model Families and Versions
Figure 13 provides a comprehensive taxonomy of Multimodal Large Language Models (MLLMs), summarizing their diversity across families, versions, and first-generation releases. The diagram branches multiple versions of a model family (e.g., InternVL, LLaVA, VILA, Phi3, Ovis) and separates earlier first-generation models (e.g., Video-LLaVA Lin et al. (2024a); Zhu et al. (2024), Vintern Doan et al. (2024), LLama3.2-Vision Dubey et al. (2024); Meta (2024)) to reflect their evolutionary development, domain-specific popularity and capabilities in conversational systems Pattnayak et al. (2025a, b, ); Patel et al. (2025), finetuning Thomas et al. (2025), and handling uses across documents Yin et al. (2024); Agarwal et al. (2025b); Panda et al. (2025), images and videos .
This taxonomy highlights the wide range of MLLMs currently available, organized as follows:
  • InternVL Family: Spanning models from Intern-VL1.1 Chen et al. (2024c); Gao et al. (2024); Chen et al. (2024b) to the advanced Intern-VL2.5-MPO Chen et al. (2024d), this family emphasizes efficiency and adaptability for diverse tampering and multimodal tasks. Successive iterations demonstrate marked improvements, particularly in handling temporal disruptions such as Dropping and Substitution. The InternVL2-5-8B models showcases its ability to handle fine-grained spatiotemporal reasoning, highlighting its dominance in high-resource benchmarks.
  • LLaVA Family: Starting with LLaVa-NEXT Li et al. (2024d), which struggled across most benchmarks, this family has evolved with models like LLaVa-OneVision Li et al. (2024b) and LLaVa-Video Zhang et al. (2024), demonstrating significant improvements in task-specific video understanding through optimized pretraining and alignment techniques. Despite the advancements, LLaVa-OneVision & LLaVa-Video continues to face challenges in handling complex temporal disruptions, unlike Chat-UniVi, which has emerged as a robust alternative.
  • VILA Family: The VILA series, including VILA-1.5 Lin et al. (2024b), LongVILA Xue et al. (2024), and VILA-U Wu et al. (2024b), demonstrates exceptional overall performance due to robust training pipelines and innovative architectures. Notably, VILA-40B excels across benchmarks, which demonstrates advanced fine-grained and long-form video understanding capabilities, particularly in addressing tampering types like Masking and Rotation. Its architectural design allows it to efficiently process long-context visual inputs, setting a new standard for performance in large-scale benchmarks.
  • Phi3 Family: Known for its scalable architecture and focus on real-world applicability, the Phi3-Vision Abdin et al. (2024) models struggle across tampering scenarios. The Phi3.5-Vision model, in particular, highlights the benefits of improved tokenization and training strategies. It outperforms it’s earlier versions by leveraging better alignment between visual and textual modalities.
  • Ovis Family: Optimized for fine-grained visual-text alignment, Ovis Lu et al. (2024) models leverage specialized datasets for tasks requiring high-resolution image interpretation and contextual reasoning. While the smaller version struggled with temporal coherence, subsequent larger model versions show promising improvements in spatial reasoning tasks like Masking and Rotate. Such enhancement results from better alignment between text and vision modalities.
  • First-Generation Releases: Models such as Chat-UniVi Jin et al. (2024), Molmo Deitke et al. (2024), NVLM Dai et al. (2024), and Pixtral Agrawal et al. (2024) represent earlier efforts in video-language modeling. While some, like Molmo and Aria Li et al. (2024c), continue to show competitive performance due to innovative training strategies, others, such as NVLM, are limited by suboptimal optimization for temporal reasoning, which hinders their ability to adapt to tampering scenarios.
  • Qwen-VL Family: The Qwen2-VL Wang et al. (2024b) and Qwen-VL Yang et al. (2024) models are recent entrants that combine advanced architectures with scalable parameterization. They achieve strong results in grounding and visual reasoning tasks but struggle in detecting tampering across task and scenarios. Scaling the model size does help the performance but is still below the average performance of models in the study.
(a) MLLMs in the Low Performing Category
(b) MLLMs in the High Performing Category
(c) MLLMs in the Moderate Performing Category
Figure 14: F1 (overall) performance of individual models across low, moderate, and high-performing categories. Models like InternVL-2.5 lead high-performing groups, while Llava-OneVision models underperform consistently.
A.4 Extended Results & Analysis
A.4.1 Analysis based on Performance Categories
Figure 14 examines individual model performance across all categories, providing insights into trends among model families, versions, and architectures.
For low-performing models (Figure 14(a)), Llava-OneVision exhibits consistently weak performance across tampering types, even at larger parameter sizes (e.g., Llava-OneVision-72B). This suggests potential architectural and training data limitations, particularly for temporal coherence tasks. Interestingly, Qwen2-VL-7B underperforms significantly compared to its larger counterpart, Qwen2-VL-72B, which achieves a notable improvement. This indicates that increasing model size, combined with its training paradigm, positively impacts robustness for this family.
Perf. Model Vision Language Alignment Strategy Training Strategy
High Aria SigLIP-400M Aria-MoE MoE decoder + lightweight ViT 4-stage: language, multimodal, long-context, post-training
High LLaMA In-House LLaMA-3.2 Iterative SFT, rejection sampling, DPO on curated data
High VILA InternViT-6B Yi-34B Deep embedding projection + joint tuning 3-stage: projector init, visual pretrain, instruction tuning
High InternVL InternViT-6B v2.5 Qwen2.5 QLLaMA middleware harmonization Contrastive → generative → supervised tuning
Moderate Vintern InternViT-300M Qwen2.5 MLP projector w/ visual instruction tuning 2-stage: full-param + LoRA, cross-entropy loss
Moderate Qwen2-VL Custom ViT Qwen2 Dynamic resolution + M-RoPE 3-stage: image–text pretrain + instruction tuning
Moderate Pixtral ViT-400M Nemo-12B Decoder fusion, ROPE-2D, sequence packing Interleaved image–text pretraining
Moderate NVLM InternViT-6B Qwen2 Decoder/cross-attn/hybrid variants Pre-align (frozen LLM), then multimodal SFT
Moderate Molmo CLIP ViT-L/14 Qwen2-72B Multi-crop connector + decoder LLM Length-conditioned captions → multitask tuning
Moderate Chat-UniVi CLIP-ViT-L/14-336 Vicuna-v1.5-7B Unified visual tokens + multi-scale semantics 2-stage: frozen pretrain + joint instruction tuning
Low PHI CLIP ViT-L/14 Phi-3 Transformer decoder (block-sparse) 2-phase: filtered web + synthetic
Table 6: Architecture, alignment, and training strategies for MLLMs stratified by robustness category.
In the moderate-performing category (Figure 14(c)), several trends emerge. Llama3.2-11B and Llama3.2-90B exhibit very similar performance despite the significant increase in size. These models, trained on the same recipe and dataset, highlight that merely increasing parameter count does not drastically enhance tampering detection capabilities. Another interesting observation is the improvement shown by Phi3.5-Vision and Vintern-Beta over their predecessors (Phi3-Vision and Vintern). This suggests that targeted architectural modifications or additional task-specific training significantly contribute to robustness.
For high-performing models (Figure 14(b)), InternVL-2.5 dominates across tampering effects, with smaller versions (e.g., 4B) achieving comparable performance to their larger counterparts. This demonstrates that efficient architectural design and training strategies can offset limitations in model size. VILA1.5-40B, a model specifically designed for long-form video understanding, showcases exceptional robustness, emphasizing the importance of task-specific optimization. Notably, Molmo-72B and NVLM-72B exhibit below-average performance relative to other large models or their smaller counter parts, indicating inefficiencies in parameter utilization or potential overfitting to pretraining data and tasks.
(a) Performance Distribution of Small(<7B) MLLMs
(b) Performance Distribution of Large (>26B) MLLMs
(c) Performance Distribution of Medium(>=7B & <=26B) MLLMs
Figure 15: F1 (overall) performance of individual models across small, medium, and large model size categories. Models like InternVL-2.5 lead high-performing groups, while Llava-OneVision models underperform across categories.
Another noteworthy observation across categories is the consistency within certain model families. For example, the Llava-Video and Chat-UniVi models outperform their Llava-OneVision counterparts, demonstrating the importance of video-specific training for tampering detection. Conversely, while Molmo-1B excels among small models, its larger variant (Molmo-72B) does not scale proportionally in performance, reinforcing the need for efficient scaling and training strategies.
A.4.2 Analysis based on Model Architectures
To better understand model performance under video tampering, we conducted a comparative analysis of MLLMs’ architectural design, alignment strategies, and training paradigms. Table 6 summarizes key configurations across models stratified by their performance category (low, moderate, high).
Our findings suggest that robust MLLMs typically adopt deeper integration strategies between vision and language modalities. High-performing models like VILA, InternVL, and Aria utilize multi-stage training pipelines, explicit visual-text alignment layers (e.g., projectors or middleware), and instruction tuning on curated or human-annotated datasets. In contrast, moderate-performing models often rely on lightweight fusion (e.g., MLPs or prompt tuning) or lack post-alignment tuning stages. Low-performing models tend to use frozen CLIP encoders and shallow decoders with minimal multimodal alignment.
These insights support the hypothesis that robustness to spatiotemporal manipulations correlates with stronger cross-modal alignment and iterative supervision across training stages.
A.4.3 Analysis based on Model Size
A closer examination of individual model trends across size categories (Figure 15) reveals several noteworthy patterns.
Small Models (<7B).
Among small models (Figure 15(a)), Molmo-1B demonstrates exceptional robustness, outperforming several medium-sized models and achieving consistency across all tampering types. Another notable small model, Phi3.5-Vision, shows drastic improvement over its predecessor, Phi3-Vision, highlighting the impact of architectural updates and extended task-specific training.
For the VILA model family, VILA1.5-3B performs better than many other small models, showcasing the benefits of targeted optimization for long-form video understanding. However, its performance lags behind Molmo-1B and Phi3.5-Vision, indicating room for improvement in handling complex tampering scenarios.
Interestingly, InternVL-2.5-4B matches or exceeds the performance of several medium-sized models, emphasizing the importance of efficient design over raw parameter counts. Models like Llava-OneVision-7B, however, remain among the weakest performers, suggesting that training approaches and architectural focus on static data limit their ability to handle tampering.
Medium Models (7B–26B).
For medium-sized models (Figure 15(c)), we observe considerable variability. Variants such as Molmo-D and Molmo-O exhibit similar performance, suggesting limited scalability within the family. Conversely, models like InternVL-2.5-8B maintain exceptional robustness, outperforming even larger models in the same category.
Pixtral, a medium-sized model that claims higher performance on other benchmarks, delivers average results here, highlighting that tampering-specific robustness requires distinct optimization strategies. Similarly, Chat-UniVi demonstrates an advantage over other Llava-family models, reinforcing the role of video-specific training in achieving higher tampering detection performance.
In the VILA model family, VILA1.5-8B underperforms compared to both smaller and larger VILA models, making it an interesting anomaly. This drop in performance could stem from suboptimal parameter tuning or an architectural bottleneck that affects scalability. Meanwhile, VILA1.5-13B and VILA1.5-40B continue to improve with increasing size, emphasizing the scalability of this family for long-form video tampering tasks.
Figure 16: F1 (overall) scores across tampering types for task categories. Masking is less disruptive, while Dropping and Substitution degrade performance in complex tasks like Counterfactual Inference.
Large Models (>26B).
Among large models (Figure 15(b)), we observe mixed performance trends. While InternVL-2.5 continues to dominate, its smaller variants (e.g., 4B, 8B) achieve comparable results, raising questions about the marginal benefits of scaling up within this family. VILA1.5-40B, specifically designed for long-form video understanding, ranks among the best-performing large models, showcasing the value of video-specific training & optimization.
The Qwen2-VL family highlights the importance of scaling when paired with architectural optimization. Qwen2-VL-72B significantly outperforms its 7B counterpart, demonstrating the advantages of increased parameter counts and more extensive pretraining. However, its performance still lags behind other large models like InternVL-2.5-40B and VILA1.5-40B, suggesting potential inefficiencies in architecture or video-specific optimization.
The Llama3.2-Vision family offers a nuanced perspective on scaling. While Llama3.2-11B and Llama3.2-90B exhibit very similar performance, underscoring that increasing parameter count alone does not yield proportional performance gains. This trend reflects the importance of complementing scaling with architectural innovations and diverse, tamper-focused training data.
While InternVL-2.5 continues to dominate, its smaller variants (e.g., 4B, 8B) achieve comparable results, raising questions about the marginal benefits of scaling up within this family. Similarly, Molmo-72B, despite its size, fails to match the performance of its smaller counterparts, indicating inefficiencies in parameter utilization or overfitting to pretraining data.
In contrast, Qwen2-VL-72B delivers subpar results relative to its size but still outperforms its smaller sibling, Qwen2-VL-7B. This highlights that while scaling can improve performance, architectural and training advancements are critical for leveraging the full potential of larger models.
A.4.4 Analysis across Video Task Types
Insights across Tampering Types.
Figure 16 reveals nuanced differences in performance across tampering types and video tasks. Masking consistently emerges as the least disruptive effect, particularly in tasks like Moving Direction and Object Existence, where models rely more on contextual cues than on fine-grained spatial details. Conversely, Dropping and Repetition create the most significant challenges, particularly in tasks involving long-term temporal dependencies, such as Action Sequence and Counterfactual Inference.
Notably, the performance trends also vary across model categories. High-performing models demonstrate consistent robustness across all tasks and tampering types, while low-performing models exhibit the highest susceptibility to temporal and spatial disruptions. Moderate-performing models, on the other hand, display a mix of strengths and weaknesses, excelling in tasks with static or localized changes but struggling with tasks requiring temporal coherence.
Trends across Task Categories.
Examining the task-wise trends, several key observations emerge:
  • Tasks involving long-term temporal reasoning (e.g., Counterfactual Inference, Action Sequence) are more sensitive to tampering, particularly Dropping, Repetition, and Substitution, which disrupt narrative continuity.
  • Tasks with localized changes (e.g., State Change, Episodic Reasoning) are less affected by tampering, as models rely on static visual cues.
  • Tasks requiring fine-grained spatial understanding (e.g., Fine-Grained Action) show significant degradation under Rotation, indicating a limitation in models’ ability to process spatial distortions.
  • Performance disparities across tampering types highlight architectural strengths and weaknesses. For instance, InternVL-2.5 excels in tasks like Action Prediction and Counterfactual Inference, owing to its advanced temporal reasoning capabilities.
A.5 Benchmarking Efforts
Our benchmarking efforts encompass 45 models across diverse categories and tampering scenarios, including Drop, Mask, Repeat, Rotate, and Substitute (Table 7). The results reveal significant variability in performance, influenced by model size, architecture, and training data.
Models such as VILA1.5-40B and InternVL2.5-8B emerged as top performers, achieving consistent resilience across all tampering types, with overall scores of 0.879 and 0.875, respectively. This highlights the importance of architectural innovations and advanced training techniques for tampering robustness. In contrast, early-generation models like LLaVA-OneVision underperformed across all categories, with overall scores as low as 0.001, reflecting limitations in temporal coherence and token alignment.
Specialized models for video tasks, such as Chat-UniVi and Video-LLaVA, demonstrated substantial improvements over base LLaVA models. Chat-UniVi-7B-v1.5 achieved an overall score of 0.658, significantly outperforming LLaVA-OneVision, showcasing its ability to handle complex temporal manipulations. Meanwhile, Video-LLaVA-7B-HF maintained robust performance across categories, further validating the effectiveness of unified tokenization and video-specific optimizations. However, LLaVA-Video, despite efforts to improve alignment and pretraining, continues to struggle with certain tampering types, reflecting the challenges of adapting image-centric architectures to video modalities.
Interestingly, medium-sized models like Phi3.5-Vision demonstrated notable performance improvements compared to earlier iterations such as Phi3-Vision, indicating that scaling alone does not account for robustness gains. Specialized models like Ovis1.6-Gemma2-9B showcased strengths in spatial tampering scenarios (e.g., Mask and Rotate) but struggled with temporal disruptions like Repeat and Substitute. This trend underscores the importance of task-specific optimizations.
Future Benchmarking Plans.
While our analysis has covered an extensive set of models, several promising entries are yet to be evaluated. Models such as NVILA Liu et al. (2024), LongVILA Xue et al. (2024), and AuroraCap Chai et al. (2024) are currently being integrated into our benchmarking framework, with active collaborations underway to ensure seamless evaluations. Early insights suggest that these models could offer competitive performance in handling long-form video tampering and multimodal reasoning.
Additionally, we plan to expand the scope of evaluations for models like InternVL-1, MovieChat Song et al. (2024a, b), Vintern, and future iterations of Chat-UniVi and Video-LLaVA. While Chat-UniVi has shown impressive robustness across temporal and spatial tampering scenarios, and Video-LLaVA continues to improve, exploring these models under additional tampering techniques will provide deeper insights into their limitations and areas for refinement.
The dynamic and evolving nature of our benchmarking framework ensures that future evaluations will continue to capture advancements in architectural design and tampering robustness.
A.6 Key Findings
Consistent Performers.
The InternVL-2.5 series consistently outperforms other models by achieving strong F1 (overall) scores across all tampering types. Notably, even smaller variants like InternVL-2.5-4B match the robustness of larger models. Such results highlight the efficiency of its architecture and training strategy.
Effect-specific Strengths.
Models such as Phi3.5-Vision excel in detecting Masking tampering, which indicates its specialized capabilities for handling visual obfuscations. Similarly, the VILA1.5-40B, designed for long-form video understanding, excels in spatial-temporal tasks.
Weaker Models.
Certain models, including Llava-OneVision variants, exhibit consistent weaknesses, particularly with temporal disruptions like Dropping and Repetition. This result may suggest limitations in their architectural designs and training paradigms.
Tampering Insights.
Dropping and Repetition emerge as the most challenging tampering types for all model categories, reflecting the difficulty of maintaining temporal coherence under such manipulations. In contrast, Masking is relatively less disruptive, particularly for tasks relying on contextual cues.
A.7 Discussion and Future Directions
The findings shed light on the importance of task-specific optimization and tamper-aware training for improving model robustness. Tasks that involve complex temporal dependencies or fine-grained spatial reasoning highlight critical gaps in current architectures, which need to better integrate temporal embeddings and multi-scale spatial attention mechanisms. MVTamperBench highlights the critical importance of robust architectures and diverse training data & strategies for achieving tampering resilience. Below, we outline actionable insights and avenues for future exploration:
Expanding Benchmark Scope.
To enhance the benchmark’s comprehensiveness, future iterations could evaluate additional model families, including emerging MLLMs. This will ensure a broader understanding of robustness trends.
Addressing Weak Models.
Models like Llava-OneVision & Qwen-2-VL highlight the need for targeted improvements. Techniques such as adversarial training, task-specific fine-tuning, and architectural enhancements could improve performance.
Introducing New Tampering Types.
Expanding the benchmark to include tampering techniques such as localized masking, noise injection, and frame-level shuffling would provide a more nuanced evaluation of model resilience. Additionally, exploring domain-specific tampering types for critical applications like healthcare and surveillance could reveal context-dependent vulnerabilities.
Task-specific Insights.
Examining performance at a finer granularity—e.g., evaluating models on specific task categories or within specialized domains—could provide actionable guidance for training and optimization.
Scaling Considerations.
Observations such as diminishing returns for models like Molmo-72B and Llama-3.2-90B-Vision emphasize the need for efficient scaling strategies. Future work could explore methods for optimizing parameter utilization and balancing architectural complexity with training data diversity.
Integration with Real-world Applications.
Extending MVTamperBench to evaluate tampering resilience in real-world domains like media verification, misinformation detection, and legal forensics could uncover application-specific challenges and provide a pathway for practical deployments.
By exploring these directions, we hope that our MVTamperBench will evolve as a cornerstone benchmark for tampering detection, and further drive innovation in tamper-resilient MLLMs and foster trust in their real-world applications.
Model Size Drop Mask Repeat Rotate Substitute Overall Performance Size
Category Category
Phi-3-Vision 4 0.002 0.002 0.000 0.002 0.002 0.001 Low Small
llava-onevision-qwen2-0.5b-ov 1 0.001 0.001 0.001 0.001 0.001 0.001 Low Small
Llama-3-VILA1.5-8b 8 0.001 0.003 0.001 0.003 0.001 0.002 Low Medium
llava_video_qwen2_7b 7 0.006 0.006 0.005 0.005 0.006 0.006 Low Medium
Qwen2-VL-7B-Instruct 7 0.005 0.004 0.004 0.025 0.005 0.009 Low Medium
Vintern-1B-v2 1 0.013 0.011 0.012 0.011 0.011 0.012 Low Small
llava-onevision-qwen2-7b-ov 7 0.016 0.038 0.010 0.019 0.016 0.020 Low Medium
llava-onevision-qwen2-7b-ov-chat 7 0.020 0.050 0.013 0.027 0.020 0.026 Low Medium
llava-onevision-qwen2-72b-ov-chat 72 0.021 0.125 0.018 0.034 0.021 0.044 Low Large
llava-onevision-qwen2-72b-ov 72 0.020 0.132 0.017 0.033 0.021 0.045 Low Large
Ovis1.6-Llama3.2-3B 3 0.057 0.064 0.050 0.053 0.058 0.057 Low Small
InternVL2-1B 1 0.074 0.070 0.067 0.074 0.074 0.072 Low Small
llava_video_qwen2_72b 72 0.019 0.254 0.017 0.053 0.022 0.073 Low Large
Qwen2-VL-72B-Instruct 72 0.279 0.544 0.282 0.374 0.284 0.352 Moderate Large
VILA1.5-13b 13 0.378 0.422 0.383 0.391 0.382 0.391 Moderate Medium
Pixtral-12B 12 0.452 0.469 0.453 0.454 0.453 0.456 Moderate Medium
VILA1.5-3b 3 0.457 0.609 0.469 0.536 0.470 0.508 Moderate Small
InternVL2_5-2B 2 0.500 0.538 0.505 0.523 0.499 0.513 Moderate Small
InternVL2_5-2B-MPO 2 0.518 0.558 0.519 0.534 0.517 0.529 Moderate Small
NVLM 72 0.588 0.595 0.598 0.595 0.588 0.593 Moderate Large
Ovis1.6-Gemma2-9B 9 0.600 0.593 0.600 0.602 0.600 0.599 Moderate Medium
Qwen2-VL-2B-Instruct 2 0.610 0.594 0.606 0.601 0.612 0.605 Moderate Small
Chat-UniVi-7B-v1.5 7 0.662 0.642 0.663 0.661 0.661 0.658 Moderate Medium
Chat-UniVi-7B 7 0.667 0.666 0.666 0.666 0.666 0.666 Moderate Medium
molmo-7B-O-0924 7 0.667 0.667 0.667 0.667 0.667 0.667 Moderate Medium
Video-LLaVA-7B-HF 7 0.667 0.667 0.667 0.667 0.667 0.667 Moderate Medium
molmo-72B-0924 72 0.667 0.667 0.667 0.667 0.667 0.667 Moderate Large
Vintern-3B-beta 3 0.670 0.669 0.674 0.669 0.670 0.670 Moderate Small
Phi-3.5-Vision 4 0.676 0.822 0.677 0.682 0.677 0.707 Moderate Small
InternVL2-8B 8 0.705 0.761 0.689 0.740 0.711 0.721 Moderate Medium
Aria 25 0.717 0.738 0.716 0.719 0.716 0.721 Moderate Medium
molmo-7B-D-0924 7 0.833 0.833 0.832 0.833 0.833 0.833 High Medium
molmoE-1B-0924 1 0.842 0.842 0.842 0.842 0.842 0.842 High Small
Llama-3.2-11B-Vision-Instruct 11 0.846 0.847 0.848 0.847 0.845 0.847 High Medium
InternVL2_5-26B-MPO 26 0.848 0.848 0.847 0.848 0.848 0.848 High Large
Llama-3.2-90B-Vision-Instruct 90 0.854 0.853 0.853 0.853 0.854 0.853 High Large
InternVL2_5-26B 26 0.858 0.858 0.858 0.858 0.858 0.858 High Large
InternVL2_5-4B 4 0.860 0.861 0.860 0.861 0.860 0.860 High Small
InternVL2_5-38B-MPO 38 0.857 0.868 0.859 0.868 0.859 0.862 High Large
InternVL2_5-8B-MPO 8 0.864 0.864 0.864 0.864 0.864 0.864 High Medium
InternVL2-40B 40 0.863 0.870 0.864 0.870 0.865 0.866 High Large
InternVL2_5-38B 38 0.871 0.875 0.872 0.875 0.871 0.873 High Large
InternVL2-26B 26 0.872 0.876 0.872 0.874 0.873 0.873 High Large
InternVL2_5-8B 8 0.875 0.875 0.875 0.875 0.875 0.875 High Medium
VILA1.5-40b 40 0.879 0.880 0.878 0.880 0.879 0.879 High Large
Table 7: Performance metrics for various models.